OpenStudy (anonymous):

If P (A) = 0.2 and P (B) = 0.3 and A and B are disjoint, what is P (A or B)? 0.00 0.06 0.10 0.44 0.50

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

p(A or B) =p(A)+p(B) -P(A and B) but here since A and B are disjoint hence we can say P(A and B) =0 so here p(A or B) =p(A)+p(B) =0.2+0.3 =0.50

OpenStudy (anonymous):

Thank you! I'll give you a medal, can you help with this one? If P (A) = 0.2 and P (B) = 0.3 and A and B are independent but NOT disjoint, find P (A or B). 0.06 0.10 0.44 0.50 Cannot be determined from the information given.

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

@tabicyrus It seems medals are more dear to u....?

OpenStudy (anonymous):

no but some people wont help unless I give a medal :(

OpenStudy (kropot72):

\[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] P(A or B) = 0.2 + 0.3 - (0.2 * 0.3) = you can calculate

OpenStudy (anonymous):

0.1?

OpenStudy (kropot72):

\[0.2+0.3-(0.2\times0.3)=?\]

OpenStudy (anonymous):

@tabicyrus Oh i see.. but you shouldn't say. that way...

OpenStudy (anonymous):

okay now i got .44

OpenStudy (kropot72):

@tabicyrus You are correct :)

OpenStudy (anonymous):

thank you! @kropot72 :)

OpenStudy (kropot72):

You're welcome :)